OpenStudy (anonymous):

Hi! Please help! 4 square root of 2 times square root of 13 Thanks

OpenStudy (anonymous):

\[4\sqrt{2}\times \sqrt{13}=4\sqrt{23}\] there is nothing more you can do

OpenStudy (anonymous):

is it 4 square root of 26?

OpenStudy (anonymous):

23?..